If r= 5x, then how many tenths does 1/2 of x have?
Answer by stanbon(75887)

If r= 5x, then how many tenths does 1/2 of x have?
x = (1/5)x
(1/2)x = (1/10)x
=====================
Cheers,
Stan H.